OpenStudy (anonymous):

What is the simplest form of the expression? √2 - √10/√2 + √10

OpenStudy (mertsj):

\[\sqrt{2}-\frac{\sqrt{10}}{\sqrt{2}}+\sqrt{10}\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

Thats not an option no..

OpenStudy (mertsj):

Have I posted the question accurately?

OpenStudy (anonymous):

a) -3/2 + 1/4 √10 b) -3/2 + 1/2 √5 c)1 -1/3 √5 d) 1 + 1/2√5

OpenStudy (anonymous):

can you write the question out

OpenStudy (anonymous):

i did

OpenStudy (superdavesuper):

\[\sqrt{a*b}=\sqrt{a}*\sqrt{b}\] and 10=2*5 so....

OpenStudy (anonymous):

ao would it be d?

OpenStudy (anonymous):

so*

OpenStudy (mertsj):

are you there @Goomblet

OpenStudy (superdavesuper):

hahahaa can u show me ur steps plz? i think i read the wrong eqn from above. sorry!

OpenStudy (anonymous):

You can model the populationof a certain city between 1954-2000 by the radical function P(x) = 55,000 √x-1945. Using this model, in which year was the poplulation of that city 165,000?

OpenStudy (superdavesuper):

165000=55000*\[\sqrt{x-1945}\] OK?